OpenStudy (anonymous):

kale and merricat are making decisions about their bank accounts. kale wants to deposit $200 as a principle amount, with an interest of 2% compounded quarterly. mericat wants to deposit $200 as the principle amount, with an interest of 4% compounded monthly.

OpenStudy (anonymous):

@ganeshie8 Please help me!

jimthompson5910 (jim_thompson5910):

what do you have so far?

OpenStudy (anonymous):

Nothing :( Well, I have a formula?

jimthompson5910 (jim_thompson5910):

which formula do you have

OpenStudy (anonymous):

a=p(1+r/n)^nt

OpenStudy (anonymous):

I could be wrong

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that formula will be used here

OpenStudy (anonymous):

Yay:)

jimthompson5910 (jim_thompson5910):

`Harrison wants to deposit $200 as a principle amount, with an interest of 2% compounded quarterly` P = 200 is the principal r = 0.02 is the interest rate in decimal form (2% = 2/100 = 0.02) n = 4 is the number of times you compound the money per year (quarterly = 4 times) t = 2 years

OpenStudy (anonymous):

ok so then we would just plug it in right? n=200(1+0.02/4)^(2)(4) right?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

then compute it with a calculator

OpenStudy (anonymous):

I got 204.0401

jimthompson5910 (jim_thompson5910):

I'm getting something slightly higher

OpenStudy (anonymous):

Really? Maybe I typed it in wrong

OpenStudy (anonymous):

oh okay I got 408.0602! Sorry :)

jimthompson5910 (jim_thompson5910):

did you type in `200*(1+0.02/4)^(4*2)` ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got 208.141409 when I did that

jimthompson5910 (jim_thompson5910):

I got that too

jimthompson5910 (jim_thompson5910):

208.141409 rounds to 208.14

OpenStudy (anonymous):

Ok :) What do we do next?

jimthompson5910 (jim_thompson5910):

now we move onto the next part `Sherrie wants to deposit $200 as the principle amount, with an interest of 4% compounded monthly.` what are P, r, n and t?

OpenStudy (anonymous):

p=200 r=0.04 t=12? bcuz 12 months in a year

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

oh wait, no n = 12 and t = 2 since the time frame is 2 years

OpenStudy (anonymous):

oh sorry :)

jimthompson5910 (jim_thompson5910):

I forgot to post that `Explain which method results in more money after 2 years.` so each person has their money in for t = 2 years

OpenStudy (anonymous):

Hahaha it's fine:) I get it now

OpenStudy (anonymous):

do we plug those value in the same formula?

jimthompson5910 (jim_thompson5910):

yes you'll plug in P = 200, r = 0.04, n = 12 and t = 2

OpenStudy (anonymous):

K hold on please

OpenStudy (anonymous):

216.63?

jimthompson5910 (jim_thompson5910):

I'm getting the same thing

OpenStudy (anonymous):

So the answer would be Sherrie?

jimthompson5910 (jim_thompson5910):

Harrison will have 208.14 Sherrie will have 216.63

jimthompson5910 (jim_thompson5910):

yep she has more money

OpenStudy (anonymous):

Thank-you so much!!!!

jimthompson5910 (jim_thompson5910):

no problem